OpenStudy (anonymous):

Solve each equation for x. Convert to an exponential equation. Simplify and solve the equation to find the solution. 5=log_3(x^2+18) 2=log(3x+4)

OpenStudy (johnweldon1993):

For number 1, We can raise each side to the power 3 \[\large 5 = log_3(x^2 + 18)\] \[\large 3^5 = 3^{log_3(x^2 + 18)}\] Now remember that \(\large 3^{(log_3(x))} = x\) so we have \[\large 3^5 = x^2 + 18\] \[\large 243 = x^2 + 18\] \[\large x^2 = 225\] \[\large x = \pm 15\] Did that make sense?

OpenStudy (anonymous):

Yeah thank you!

OpenStudy (johnweldon1993):

No problem, and you can use basically the same thing for question 2, however if there is no base of the log, we use the power 10

OpenStudy (anonymous):

Kkay and for the first one 3^(log_3(x)) = x is the exponential equation right

OpenStudy (johnweldon1993):

Well THAT is a way of simplifying logarithms. The actual exponential equation came from that \[\large 3^5 = 3^{log_3(x^2 + 18)}\]

OpenStudy (anonymous):

Ohh alright thanks

OpenStudy (anonymous):

could you help me on another one? I'm about to post it..

OpenStudy (johnweldon1993):

Sure :)